OpenStudy (anonymous):

The angles shown below are supplementary: A pair of supplementary angles is shown. One angle measures 10x and the other angle measures 70 degrees. What is the value of x? 10° 11° 20° 21°

OpenStudy (dayakar):

11

OpenStudy (dayakar):

10x+70=180 10x=180-70 10x= 110 x=110/10 x=11